# Diffusion Attribution Score: Evaluating Training Data Influence in Diffusion Model

### v2

@@ F.3 Baselines
Two works focus on diffusion model that also fall outside our framework. Dai & Gifford (2023) propose a method for training data attribution on diffusion models using machine unlearning (Bourtoule et al., 2021); however, their approach necessitates a specific machine unlearning training process, making it non-post-hoc and thus unsuitable for standard settings. Similarly, Wang et al. (2023) acknowledge the current challenges in conducting post-hoc training influence analysis with existing methods. They suggest an alternative termed ”customization,” which involves adapting or tuning a pretrained text-to-image model through a specially designed training procedure.
@@
where $\lambda{\bm{I}}$ is included for numerical stability and regularization. The impact of this term is further explored in Appendix G.
D-TRAK. Simliar to TRAK, as elaborated in Section 2, we adapt the D-TRAK (Zheng et al., 2024) as detailed in Eq 6. We implent the model output function $f({\bm{z}},\theta)$ as $\mathcal{L}_{\text{Square}}$. The D-TRAK is implemented using the following equations:
@@ F.4 Experiments Result
Appendix G Ablation Studies
We conduct additional ablation studies to evaluate the performance differences between D-TRAK and DAS. In this section, CIFAR-2 serves as our primary setting. Further details on these settings are available in Appendix F.1. We establish the corresponding LDS benchmarks as outlined in Appendix F.2.
@@ Appendix G Ablation Studies
Value of $\lambda$ In our computation of the inverse of the Hessian matrix within the DAS framework, we incorporate the regularization parameter $\lambda$, as recommended by Hastie (2020), to ensure numerical stability and effective regularization. Traditionally, $\lambda$ is set to a value close to zero; however, in our experiments, a larger $\lambda$ proved necessary, possibly due to the absence of normalization by the number of training samples $N$ in the kernel (Park et al., 2023).
@@
(c) VAl-1000
(d) GEN-10
(e) GEN-100
(f) GEN-1000
Appendix H Limitations and Broader Impacts
H.1 Limitations

### v4

@@ E.3 Baselines
Two works focus on diffusion model that also fall outside our framework. Dai & Gifford (2023) propose a method for training data attribution on diffusion models using machine unlearning; however, their approach necessitates a specific machine unlearning training process, making it non-post-hoc and thus unsuitable for standard settings. Similarly, Wang et al. (2023) acknowledge the current challenges in conducting post-hoc training influence analysis with existing methods. They suggest an alternative termed ”customization,” which involves adapting or tuning a pretrained text-to-image model through a specially designed training procedure.
(a) VAL-10
(b) VAL-100
(c) GEN-10
(d) GEN-100
Figure 5: LDS (%) on ArtBench-2 under different $\lambda$. We consider 10 and 100 timesteps selected to be evenly spaced within the interval $[1,T]$, which are used to approximate the expectation $\mathbb{E}_{t}$. We set $k=32768$.
@@
(a) VAL-10
(b) VAL-100
(c) GEN-10
(d) GEN-100
Figure 6: LDS (%) on ArtBench-5 under different $\lambda$. We consider 10 and 100 timesteps selected to be evenly spaced within the interval $[1,T]$, which are used to approximate the expectation $\mathbb{E}_{t}$. We set $k=32768$.
@@
Table 3: LDS (%) on CelebA with timesteps (10 or 100)
Results on CelebA
Method Validation Generation
10 100 10 100
Raw pixel (dot prod.) 5.58$\pm$0.73 -4.94$\pm$1.58
Raw pixel (cosine) 6.16$\pm$0.75 -4.38$\pm$1.63
CLIP similarity (dot prod.) 8.87$\pm$1.14 2.51$\pm$1.13
CLIP similarity (cosine) 10.92$\pm$0.87 3.03$\pm$1.13
TracInCP 5.14$\pm$0.75 4.89$\pm$0.86 5.18$\pm$1.05 4.50$\pm$0.93
GAS 5.44$\pm$0.68 5.19$\pm$0.64 4.69$\pm$0.97 3.98$\pm$0.97
Journey TRAK / / 6.53$\pm$1.06 10.87$\pm$0.84
Relative IF 11.10$\pm$0.51 19.89$\pm$0.50 6.80$\pm$0.77 14.66$\pm$0.70
Renorm. IF 11.01$\pm$0.50 18.67$\pm$0.51 6.74$\pm$0.82 13.24$\pm$0.71
TRAK 11.28$\pm$0.47 20.02$\pm$0.47 7.02$\pm$0.89 14.71$\pm$0.70
D-TRAK 22.83$\pm$0.51 28.69$\pm$0.44 16.84$\pm$0.54 21.47$\pm$0.48
DAS 29.38$\pm$0.51 33.79$\pm$0.23 28.73$\pm$0.49 30.68$\pm$0.31
@@
where $\lambda{\bm{I}}$ is included for numerical stability and regularization. The impact of this term is further explored in Appendix F.
D-TRAK. Similar to TRAK, as elaborated in Section 3, we adapt the D-TRAK (Zheng et al., 2024) as detailed in Eq 6. We implement the model output function $f({\bm{z}},\theta)$ as $\mathcal{L}_{\text{Square}}$. The D-TRAK is implemented using the following equations:
@@ E.4 Experiments Result
Appendix F Ablation Studies
We conduct additional ablation studies to evaluate the performance differences between D-TRAK and DAS. In this section, CIFAR-2 serves as our primary setting. Further details on these settings are available in Appendix E.1. We establish the corresponding LDS benchmarks as outlined in Appendix E.2.
@@ Appendix F Ablation Studies
Value of $\lambda$ In our computation of the inverse of the Hessian matrix within the DAS framework, we incorporate the regularization parameter $\lambda$, as recommended by Hastie (2020), to ensure numerical stability and effective regularization. Traditionally, $\lambda$ is set to a value close to zero; however, in our experiments, a larger $\lambda$ proved necessary. This is because we use the generalized Gauss-Newton (GGN) matrix to approximate the Hessian in the computation of DAS. Unlike the Hessian, the GGN is positive semi-definite (PSD), meaning it does not model negative curvature in any direction. The main issue with negative curvature is that the quadratic model predicts unbounded improvement in the objective when moving in those directions. Without certain techniques, minimizing the quadratic model results in infinitely large updates along these directions. To address this, several methods have been proposed, such as the damping technique discussed in (Martens, 2020). In our paper, we adopt the linear damping technique $\lambda I$ used in (Zheng et al., 2024; Georgiev et al., 2023), which has proven effective on diffusion models. We show how $\lambda$ influence the LDS result in Figure 4, Figure 5, Figure 6, Figure 7 and Figure 8.
@@
(c) GEN-10
(d) GEN-100
Figure 7: LDS (%) on CIFAR-10 under different $\lambda$. We consider 10 and 100 timesteps selected to be evenly spaced within the interval $[1,T]$, which are used to approximate the expectation $\mathbb{E}_{t}$. We set $k=32768$.
(a) VAL-10
(b) VAL-100
(c) GEN-10
(d) GEN-100
Figure 8: LDS (%) on CelebA under different $\lambda$. We consider 10 and 100 timesteps selected to be evenly spaced within the interval $[1,T]$, which are used to approximate the expectation $\mathbb{E}_{t}$. We set $k=32768$.
(a) ArtBench2-CLIP
(b) CIFAR2-CLIP
(c) ArtBench2-$L^{2}$ Dist.
(d) CIFAR2-$L^{2}$ Dist.
Figure 9: Box-plots of counterfactual evaluation on CIFAR2 and ArtBench2. We assess the impact of removing the 1,000 highest-scoring training samples and retraining the model using Random, TRAK, D-TRAK, and DAS. The evaluation metrics include pixel-wise $L^{2}$-Distance and CLIP cosine similarity between 60 generated samples and the corresponding images generated by the retrained models, sampled from the same random seed.
Appendix G Counter Factual Experiment
Hu et al. (2024) discuss some limitations of LDS evaluation in data attribution. To further validate the effectiveness of DAS, we also conduct an counter-factual experiment, that 60 generate images are attributed by different attribution method, including TRAK, D-TRAK and DAS. We detect the top-1000 positive influencers identified by these methods and remove them from the training set and re-train the model. We utilize 100 timesteps and a projection dimension of $k=32768$ to identify the top-1000 influencers for TRAK, D-TRAK and DAS. Additionally, we conduct a baseline experiment where 1000 training images are randomly removed before retraining. The experiment is conducted on ArtBench2 and CIFAR2. We generate the new images with same random seeds and compute the pixel-wise $L^{2}$-Distance and CLIP cosine similarity between the re-generated images and their corresponding origin images. The result is reported in Figure 9 with boxplot. For the pixel-wise $L^{2}$-Distance, D-TRAK yields values of 8.97 and 187.61 for CIFAR-2 and ArtBench-2, respectively, compared to TRAK’s values of 5.90 and 168.37, while DAS results in values of 10.58 and 203.76. DAS achieves median similarities of 0.83 and 0.71 for ArtBench-2 and CIFAR-2, respectively, which are notably lower than TRAK’s values of 0.94 and 0.84, as well as D-TRAK’s values of 0.88 and 0.77, demonstrating the effectiveness of our method.
(a) Origin
(b) Counter Factual
Figure 10: The result of Toy Experiment in Appendix H. Scatter plot showing the relationship between $L^{2}$ distance and two metrics: loss difference and noise predictor output difference over entire generation. The noise predictor output difference exhibits a stronger correlation with $L^{2}$ distance, indicating its effectiveness in capturing image variation.
(a) Origin
(b) Counter Factual
(c) Origin
(d) Counter Factual
Figure 11: Figures 11(a) and 11(b) represent a pair of generated images, while 11(c) and 11(d) form another pair. Despite the loss difference between 11(a) and 11(b) being only 0.0007, the $L^{2}$ distance is 15.261. Similarly, the loss difference between 11(c) and 11(d) is also 0.0007, yet the $L^{2}$distance is 14.485, showing that the loss value fail to trace the change on generated images.
Appendix H Evaluating Output Function Effectiveness
Here, we present a toy experiment to validate our theoretical claims: using the Simple Loss Value to represent changes in generated images is inadequate, as it relies on an indirect distributional comparison, as discussed in Eq. 8. Instead, we propose using changes in the noise predictor output of the diffusion model. In the toy experiment, we first train a unconditional DDPM on CIFAR2. The origin and each retrained model are to generate an image pair where one random seed for a re-trained model. In total, we have 60 different generated images pairs. The $L^{2}$ distance between the generated and original images is calculated to directly measure their differences.
The original images are first noised to timestep $T$, and the two models are then used to denoise the latent variables at $T$. At the entire denoising process, we compute the average loss difference and the average noise predictor output difference. For the 60 pairs of generated samples, we calculate the rank correlation between these differences and the $L^{2}$ distance. The results reveal that the Pearson correlation between the $L^{2}$ distance and the average loss difference is only 0.257, while the correlation with the average noise predictor output difference reaches 0.485. This indicates that the noise predictor output difference aligns better with the $L^{2}$ distance than the loss value difference. Larger noise predictor output differences correspond to larger $L^{2}$ distances, reflecting greater image variation. A scatter plot of these results is shown in Figure 10.
Figure 11 provides two specific examples. For images 11(a) and 11(b), as well as 11(c) and 11(d), the loss differences between the models are only 0.0007, the smallest among all pairs. However, the $L^{2}$ distance between 11(a) and 11(b) is 15.261, and between 11(c) and 11(d) is 14.485, making them among the top five pairs with the largest $L^{2}$ distances in the toy dataset, which means the loss difference fails to measure the difference on images. Meanwhile, the noise output differences for these two pairs align closely with the trend line in Figure 10(b), showing a correlation with the $L^{2}$ distance.
This result is unsurprising. Consider an extreme scenario: a model trained on a dataset containing cats and dogs generates an image of a cat. If all cat samples are removed, and the model is retrained, the new model would likely generate a dog image under the same random seed. In such a case, the loss values for both images might both remain $v$, resulting in a loss difference of 0. This may occurs because the simple loss reflects the model’s convergence for the generation, not the image itself. Thus, the shift on loss value fail to capture the extent of image changes. By contrast, analyzing the differences in the noise predictor’s outputs at each timestep allows us to effectively trace the diffusion model output on the images.
Appendix I Application
Recent research has underscored the effectiveness of data attribution methods in a variety of applications. These include explaining model predictions (Koh & Liang, 2017; Ilyas et al., 2022), debugging model behaviors (Shah et al., 2023), assessing the contributions of training data (Ghorbani & Zou, 2019; Jia et al., 2019), identifying poisoned or mislabeled data (Lin et al., 2022), most influential subset selection (Hu et al., 2024) and managing data curation (Khanna et al., 2019; Liu et al., 2021; Jia et al., 2021). Additionally, the adoption of diffusion models in creative industries, as exemplified by Stable Diffusion and its variants, has grown significantly (Rombach et al., 2022; Zhang et al., 2023). This trend highlights the critical need for fair attribution methods that appropriately acknowledge and compensate artists whose works are utilized in training these models. Such methods are also crucial for addressing related legal and privacy concerns (Carlini et al., 2023; Somepalli et al., 2023).
Appendix J Limitations and Broader Impacts
J.1 Limitations